OpenStudy (anonymous):

Find the midpoint of PQ if P(2, 7) and Q(10, 13)? Answers: (10, 6) (6, 10) (-10, -6) (-6, -10)

OpenStudy (anonymous):

@ghazi

OpenStudy (ghazi):

10+2/2 =6

OpenStudy (ghazi):

13+7/2 =10 so your point is (6,10)

OpenStudy (anonymous):

thank you can you help me with another one?

OpenStudy (ghazi):

\[\frac{ x1+x2 }{ 2 }, \frac{ y1+y2 }{ 2 }\]

OpenStudy (anonymous):

yea I know the formula but Im having trouble solving

OpenStudy (anonymous):

Find the midpoint of PQ if P(-5, -2) and Q(-5, 5)?

OpenStudy (ghazi):

(-5+(-5))/2 = -5-5/2=-10/2=-5 and for y its -2+5/2= -3/2

OpenStudy (anonymous):

5 -3/2?